# thanks: M. A. C. A., N.-Y. L. and C.-H. M. contributed equally in the manuscriptthanks: M. A. C. A., N.-Y. L. and C.-H. M. contributed equally in the manuscriptthanks: M. A. C. A., N.-Y. L. and C.-H. M. contributed equally in the manuscript

### v2

@@ Abstract
The findings in Figs. 5 imply that flux crosstalk compensation creates a magnetic flux crosstalk-free digital twin of the coupler-mediated CZ gate based on empirical adjustments of the Rabi oscillation model [6, 49, 54]. This digital twin helps identify non-idealities in two-qubit gates such as reduced 2Q Rabi oscillation cycles brought upon by an effective flux transient [6, 31, 49] and unintended hybridization of the qubit and coupler modes. Shaping the flux pulse to account for flux transients [6, 31, 49] while avoiding frequencies that invoke unwanted hybridization of modes [52] could help to match the digital twin with experimental results. We consider the proposed measurement, modeling, and analysis protocol to be suitable for scaling to hundreds of qubits with the aid of high-performance computing [2] (see Sec. SVII of [33] for supporting details).
@@
In this work, we develop the MZLC protocol for characterizing flux crosstalk in a TC architecture. The method exploits residual inductive coupling between several flux-tunable elements to resolve the 0–1 transition frequencies of the probed element. Whereas other flux crosstalk characterization strategies use Rabi spectroscopy  [23, 27] and Ramsey sequences on a single-biased element to characterize the flux crosstalk matrix, relying on the readout fidelity of the qubit/coupler. MZLC has the advantage of characterizing flux crosstalk even in modest readout SNR and an arbitrary idle flux bias range. The method requires knowledge of the dependence of the qubit/coupler spectrum on flux bias, which is routine in the bring-up of superconducting processors, making the MZLC a simple, time-efficient tool for characterizing the flux crosstalk matrix of a transmon-coupler subsystem and optimizing CZ gate operation points. Lastly, compared with the CZ SWAP model, the experimental result shows a nearly identical map with zero crosstalk. Therefore, it provides a way to create a digital twin for the TC system CZ optimization.

### v3

@@ Abstract
The findings in Figs. 5 imply that flux crosstalk compensation creates a magnetic flux crosstalk-free intuitive digital twin of the coupler-mediated CZ gate. This twin is based on empirical adjustments of the Rabi oscillation model [hellings_2025], whose analytical form is validated by carefully calibrated two-qubit state measurements [Krizan_2025] and initially modelled as three-interacting qutrits [Sung_2021]. This digital twin helps identify non-idealities in two-qubit gates such as reduced 2Q Rabi oscillation cycles brought upon by an effective flux transient [Sung_2021, TMLi2024, hellings_2025] and unintended hybridization of the qubit and coupler modes. Shaping the flux pulse to account for flux transients [Sung_2021, TMLi2024, hellings_2025] while avoiding frequencies that invoke unwanted hybridization of modes [chen2024_reset] could help to match the digital twin with experimental results. We consider the proposed measurement, modeling, and analysis protocol to be suitable for scaling to hundreds of qubits with the aid of high-performance computing [Mohseni_2025] (see Sec. SVII of [SupMat] for supporting details).
@@
In this work, we develop the MZLC protocol for characterizing flux crosstalk in a TC architecture. The method exploits residual inductive coupling between several flux-tunable elements to resolve the 0–1 transition frequencies of the probed element. Whereas other flux crosstalk characterization strategies use Rabi spectroscopy  [Barrett2023, Kosen_2024] and Ramsey sequences on a single-biased element to characterize the flux crosstalk matrix, both rely on good readout fidelity of the qubit/coupler. MZLC has the advantage of characterizing flux crosstalk even in modest readout SNR and an arbitrary idle flux bias range. The method requires knowledge of the dependence of the qubit (coupler) spectrum on flux bias, which is routine in the bring-up of superconducting processors. As discussed in the supplemental material [SupMat], its primary advantage stems from avoiding the prerequisite calibration steps for precise $\pi$ and $\pi/2$ gate operations when calibrating flux crosstalk. This combination makes the MZLC a simple, scalable, robust and low-overhead tool for characterizing the flux crosstalk matrix of a transmon-coupler subsystem and optimizing CZ gate operation points. Lastly, compared with the CZ SWAP model, the experimental result shows a nearly identical map with zero crosstalk. Therefore, it provides a way to create a digital twin for the TC system CZ optimization.
Supplemental Material
The supplemental material [SupMat] includes the following: I) details of the QPU fabrication and design; II) experimental details and essential QPU parameters; III) calibration between flux and qubit-coupler interactions; IV) flux-induced inhomogeneous effects on linewidth in MZLC; V) spectroscopic signal-to-noise ratio and frequency-shift uncertainty of MZLC; VI) effect of flux crosstalk compensation on single-qubit gate fidelities; VII) scalability of flux crosstalk characterization and compensation for large-scale superconducting processors.
Data Availability
The data that support the most of the findings of this study are openly available in Zenodo at http://doi.org/10.5281/zenodo.17639715. Check the latest version for the updated datasets. All other data are available from the corresponding authors upon request.
@@ Acknowledgment
Author Declaration
Conflict of Interest
The authors have no conflicts to declare.
Author Contributions